OpenStudy (anonymous):

If tan=-3/4, sin<0, cos=?

OpenStudy (dumbcow):

oh haha you had a typo huh?

OpenStudy (anonymous):

yeah sorry haha

OpenStudy (dumbcow):

ok set up a triangle in 4th quadrant  \[\sin \theta = \frac{-3}{r}\]

OpenStudy (anonymous):

thank you thank you

OpenStudy (dumbcow):

your welcome

OpenStudy (dumbcow):

drawing the triangles and using SOH CAH TOA can really help make these problems easy

OpenStudy (anonymous):

right so cos would just be 4/r?

OpenStudy (dumbcow):

yep

OpenStudy (anonymous):

thanks a lot!